Two 20.0‑g ice cubes at −17.0 °C are placed into 225 g of water at 25.0 °C. Assuming no energy is transferred to or from the surroundings, calculate the final temperature of the water after all the ice melts. heat capacity of ?2?(?) 37.7 J/(mol⋅K) heat capacity of ?2?(?) 75.3 J/(mol⋅K) enthalpy of fusion of ?2? 6.01 kJ/mol
Answer #1

given

mass of ice cubes = 20 g

temperature of ice T1= -17

mass of water = 225 g

temperature of water T2 = 25

Let final temperature be T. Then

for ice to become water at temperature T first it should be converted to ice at 0 degrees and then water at 0 degrees and then water at T degrees.

The heat energy required to convert ice at -17 degrees to ice at 0 degrees H1 =

moles of ice* heat capacity of ice * ( final temperature - initial temperature)

=(20/18)*37.7*(0-(-17)) = 12818/18 =712 joules

Heat energy required to convert ice at 0 degrees to water at zero degrees H2 =mass of ice* enthalpy of fusion of ice

=(20/18)*6010 = 6677 j

heat energy to make ice into water at T degrees is H3

=moles of ice*heat capacity of water*(final temperature - initial temperature)

=(20/18)*75.3*(T-0)

Heat energy given by water H4 = moles of water*heat capacity of water*(initial temperature - final temperature )

= 225/18 * 75.3 *(25-T)

As no heat is from outside,

H4 = H1 + H2 +H3 +H4

23531- 941.25*T = 712 +6677 + 83.667T

T=16142/1024 = 15.76 degrees

HIT A LIKE THERE PLEASE!!!!!!!!
Answer #2
  1. Given Data:

    • Mass = 225 g

    • Initial temperature = 25.0 °C

    • Heat capacity of water,

    • Moles of water = 225g18.015g/mol=12.49mol

    • Mass = 2 × 20.0 g = 40.0 g

    • Initial temperature = −17.0 °C

    • Heat capacity of ice,

    • Molar mass of ice (H₂O) = 18.015 g/mol → Moles of ice = 40.0g18.015g/mol=2.22mol

    • Ice cubes:

    • Water:

    • Enthalpy of fusion of ice, ΔHfusion=6.01kJ/mol=6010J/mol

  2. Energy Required to Heat Ice to 0 °C:

    q1=nice×Cice×ΔT=2.22mol×37.7

  3. Energy Required to Melt Ice:

    q2=nice×ΔHfusion=2.22mol×6010J/mol=13,342J

  4. Total Energy Absorbed by Ice (qabsorbed):

    qabsorbed=q1+q2=1422J+13,342J=14,764J

  5. Energy Released by Water to Reach Final Temperature (qreleased):

    qreleased=nwater×Cwater×(25.0°CTf)14,764J=12.49mol×75.3

  6. Solve for Final Temperature (Tf):

    25.0Tf=14,76412.49×75.325.0Tf=15.7°CTf=25.0°C15.7°C=9.3°C

Answer:

The final temperature of the water after all the ice melts is 9.3 °C


answered by: anonymous